# Comparing Neighbors Together Makes it Easy: Jointly Comparing Multiple Candidates for Efficient and Effective Retrieval

### v1

@@ Abstract
A common retrieve-and-rerank paradigm involves retrieving a broad set of relevant candidates using a scalable bi-encoder, followed by expensive but more accurate cross-encoders to a limited candidate set. However, this small subset often leads to error propagation from the bi-encoders, thereby restricting the performance of the overall pipeline. To address these issues, we propose the Comparing Multiple Candidates (CMC) framework, which compares a query and multiple candidate embeddings jointly through shallow self-attention layers. While providing contextualized representations, CMC is scalable enough to handle multiple comparisons simultaneously, where comparing 2K candidates takes only twice as long as comparing 100. Practitioners can use CMC as a lightweight and effective reranker to improve top-1 accuracy. Moreover, when integrated with another retriever, CMC reranking can function as a virtually enhanced retriever. This configuration adds only negligible latency compared to using a single retriever (virtual), while significantly improving recall at K (enhanced). Through experiments, we demonstrate that CMC  as a virtually enhanced retriever, significantly improves Recall@k (+6.7, +3.5%-p for R@16, R@64) compared to the initial retrieval stage on the ZeSHEL dataset. Meanwhile, we conduct experiments for direct reranking on entity, passage, and dialogue ranking. The results indicate that CMC is not only faster (11x) than cross-encoders but also often more effective, with improved prediction performance in Wikipedia entity linking (+0.7%-p) and DSTC7 dialogue ranking (+3.3%-p). The code and link to datasets are available at https://anonymous.4open.science/r/cmd/
@@ 1 Introduction
Figure 1: Overview of ranking model architectures for sequence retrieval tasks. Model (a), (b), and (c) are existing architectures. Model (d) is our proposed ‘Comparing Multiple Candidates (CMC)’ architecture, which computes compatibility scores by comparing the embeddings of a query and K multiple candidates via self-attention layers. Contarty to (a)-(c), our model can process multiple documents at once rather than conducting multiple forward passes of (query, candidate) pair for each document.
@@
The two-stage approach of retrieval and reranking has become a predominant approach for open-domain question answering (ODQA) Nogueira and Cho (2019); Agarwal et al. (2022b); Shen et al. (2022); Qu et al. (2020), entity linking (EL) Wu et al. (2020); Zhang and Stratos (2021); Xu et al. (2023), and dialogue systems Mele et al. (2020). Typically, bi-encoders (BE) are used to efficiently retrieve relevant candidates among a large set of documents (e.g. knowledge base), and then cross-encoders (CE) effectively rerank only a confident subset of candidates already retrieved by BE Nogueira and Cho (2019) (Figure 1.a-b).
The two-stage approach, although widely used, has limitations: it has an innate trade-off between efficiency and effectiveness and is susceptible to error propagation. When inaccurate BE retrieve too few candidates, the framework risks missing the gold candidates. Once the ground truth is not retrieved, no subsequent reranking can consider the gold candidates anymore, leading to the failure of the entire system. In contrast, simply increasing the number of candidates introduces another problem of slow processing time that limits scalability under time constraints. Furthermore, increasing the number of candidates for CE does not necessarily improve end-to-end accuracy  [^1] We also confirm this in the experiments. See appendix E.6. Wu et al. (2020). Therefore, the retrieval framework poses a dilemma for users between the inaccurate performance of BE and the slow runtime of CE.
@@
To resolve this problem, several studies have explored various strategies to find an optimal balance in the efficiency-effectiveness tradeoff. Khattab and Zaharia (2020); Zhang and Stratos (2021); Cao et al. (2020); Humeau et al. (2019) have enhanced bi-encoder architectures by adding an interaction component between the query and the candidate. However, these late interaction models require to save multiple vector embeddings for each candidate in advance, or only utilize single query-candidate interaction without considering the relationships between candidates (Figrue 1.c).
@@
In this paper, we present a novel reranking framework CMC, which stands for Comparing Multiple Candidates (Figure 1.d). CMC is designed to improve both accuracy and scalability by comparing multiple candidates along with the query simultaneously. It compares a query and multiple pre-computed candidate embeddings jointly through a few layers of bi-directional self-attention layers. In other words, CMC only takes one forward pass for single input $(\text{query},\text{candidate}_{1},...,\text{candidate}_{k})$, while CE and other late interaction models take $k$ separate forward passes for multiple input pairs $(\text{query},\text{candidate}_{1}),...,(\text{query},\text{candidate}_{k})$. The proposed configuration of CMC  maintains both the efficiency of BE by using pre-computed candidate embeddings, and the effectiveness of CE by using attention mechanisms between query and multiple candidates (Figure 1.d).
Practitioners can use CMC as a fast and effective reranker, capable of enhancing top-1 retrieval (↑R@1). Also, its efficiency enables CMC to function as a virtually enhanced retrieval system. When integrated with another retriever, such as BE, CMC efficiently identifies better candidates (enhanced; ↑R@K) from a large pool with minimal additional latency (vitrual; Figure 2). As slow cross-encoders can only process a limited number of candidates, filtering higher-quality candidates with CMC contributes to minimizing the error propagation in the retrieval process.
@@
In experiments, we evaluate CMC on Zero-SHot Entity-Linking dataset (ZeSHEL; Logeswaran et al. (2019)) to investigate the performance as a virtually enhanced retriever. The results show that CMC provides higher recall than existing state-of-the-art retrievers (+3.5 Recall@64 points) at marginal computing costs (+0.07x runtime of BE; Table 5). Additionally, we demonstrate that CMC provides high-quality candidates that enhance CE’s end-to-end prediction accuracy even with a reduced candidate set, while improving overall latency (Table 2). To examine the effectiveness of CMC as a reranker, we also evaluate CMC on entity, passage, and dialogue ranking tasks. We observe that CMC outperforms CE on both Wikipedia entity linking datasets (+0.7 accuracy points) and DSTC7 dialogue ranking datasets (+3.3 MRR points), requiring only 0.09x of CE’s reranking latency (Table 3).
The main contribution of the paper is as follows:
  • We propose a novel CMC that is both effective and scalable reranker by utilizing single pre-computed candidate embeddings (§3).
  • We show that CMC can function as a virtually enhanced retriever, which increases the recall of the initial retriever at a marginal cost and thus improves overall reranking performance. (§4.3)
@@
  • We provide experimental results which show CMC has robust performance on passage ranking, entity linking, and dialogue ranking tasks compared to various baselines (§4.4).
@@
  • Additionally, we show that CMC can benefit from domain transfer from sentence encoders while BE cannot (§4.5).
@@
Figure 2: Overview of the proposed CMC framework that compares multiple candidates at once. The configuration of Bi-encoder + CMC can be used as both a virtually enhanced retriever which finds top-K’ candidates or a reranker which outputs top-1 candidate. Candidate embeddings for bi-encoders and CMC are both precomputed while query embeddings for bi-encoders and CMC are computed in parallel on the fly. After bi-encoders retrieve top-$K$ candidates, CMC indexes the corresponding candidate embeddings and passes through a two-layer transformer encoder.
@@ 2.1 Retrieve and rerank
Information retrieval systems commonly comprise two stages: efficient retrieval and effective reranking. A fast retriever will score the query $q$ with each candidate $c\in\mathcal{C}$. Although the retriever is designed to be fast, its top-1 accuracy tends to be suboptimal. Practitioners therefore identify a candidate set.
$C_{q}=\{c_{q,1},c_{q,2},\ldots,c_{q,K}\}\subseteq\mathcal{C},$ (1)
whose elements are $K$ most relevant candidates in corpus $\mathcal{C}$, to be reranked.
A reranker $s_{\theta}(q,c_{q,j})(1\leq j\leq K)$ is a model learned to assign more fine-grained matching scores between a query $q$ and each candidate $c_{q,j}$ from the relatively small set of candidates $C_{q}$ predicted by the retriever. It is an expressive model that is generally slower but more accurate than the retriever. The candidate with the highest score,
$\hat{c}_{q}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}s_{\theta}(q,c_{q,j}),$ (2)
is the final output where query $q$ should be linked.
@@ 2.2 Related Work
In the two-stage retrieval system, the compatibility score between the query and candidate can be computed by diverse functions. Nogueira et al. (2019a) first retrieves candidates for a query using the bag-of-words BM25 retriever and then uses a cross-encoders, transformer encoders that input the query and candidate tokens jointly, to get a final score. Numerous works Lee et al. (2019); Gillick et al. (2019); Karpukhin et al. (2020) use a pre-trained language model for bi-encoders architecture to encode a query and a candidate separately and get the compatibility score between the them. The scalability of bi-encoders comes from the indexing of candidates and maximum inner-product search (MIPS); however, they are less effective than cross-encoders as candidate representation does not reflect query information (Figure 1 (a-b)). To enhance the performance of bi-encoders, follow-up works propose a task-specific fine-tuned model Gao and Callan (2022), injecting graph information Wu et al. (2023); Agarwal et al. (2022a), and multi-view text representation Ma et al. (2021); Liu et al. (2023).
@@
Late interaction models, which typically function as either a retriever or a reranker, enhance bi-encoder architectures with an interaction component between the query and candidates.
Poly-encoder Humeau et al. (2019) and MixEncoder Yang et al. (2023) represent query information through cross-attention with individual candidates to calculate matching scores. However, these models have overlooked the opportunity to explore the interaction across candidates.
Sum-of-Max Khattab and Zaharia (2020); Zhang and Stratos (2021); Santhanam et al. (2021) and DeFormer Cao et al. (2020) rely on Maximum Similarity (MaxSim) operations or additional cross-encoder layers on top of bi-encoders. These approaches increase offline indexing costs as they require every token embedding for each candidate, complicating their application to tasks where the knowledge base is frequently updated and the index must be continuously revised.
CMC diverges from these models by using a single vector embedding for each candidate, enabling interactions among multiple candidates with enhanced scalability. It helps to explore deeper relational dynamics among candidates while enhancing memory efficiency.
@@
CMC is not the first approach to jointly compare a query and documents to enhance ranking performance. This listwise ranking method has been explored in previous works Han et al. (2020); Zhang et al. (2022); Xu et al. (2023). They process cross-encoder logits for the list $(\text{query},\text{candidate}_{1},\ldots,\text{candidate}_{\text{K}})$ to rerank $K$ candidates from cross-encoders. Although they focus on performance, these approaches lack scalability as depending on pairwise representations from cross-encoders.
Unlike previous listwise ranking models with pairwise representations from cross-encoders, we propose a method that employs single representations from sentence encoders. In particular, CMC can function as a virtually enhanced retriever, boosting scalability by using independent sentence encoders for query and candidate, while still maintaining accurate predictions.
@@ 3.1 Model Architecture
Comparing Multiple Candidates, CMC, employs self-attention layers to capture both query-candidate and candidate-candidate interactions. Contrary to other late interaction models Khattab and Zaharia (2020); Humeau et al. (2019); Yang et al. (2023), which compute the matching scores by only comparing a query and an individual candidate, CMC  considers the query in conjunction with all candidates simultaneously. In other words, CMC enriches each candidate’s representation by contextualizing it with other candidates and the query. In addition, this architecture enables the processing of multiple candidates in a scalable manner by pre-computing individual candidate embeddings as discussed in §3.2. The overall architecture in Figure 1.(d) provides a high-level picture of the CMC, describing how it integrates and processes multiple candidate interactions given a single query. In detail, the self-attention layer is designed to process the aggregated encoder output (i.e. $[CLS]$ token embedding) of the query and multiple candidates. The outputs are derived from our query encoder $\texttt{Enc}_{qry}$ and candidate encoder $\texttt{Enc}_{can}$.
@@
Prior to CMC, the initial retriever identifies the candidate set with K elements $C_{q}=\{c_{q,1},...c_{q,K}\}$ for query $q$. Each query and candidate is represented by tokens $\text{x}_{q}$ and $\text{x}_{c_{q,j}}$. Initially, CMC obtains the aggregated output of query sentence tokens ($\mathbf{h}^{sent}_{q}$) and candidate sentence tokens ($\mathbf{h}^{sent}_{c_{q,j}}$) from the query encoder $\texttt{Enc}_{qry}$ and the candidate encoder $\texttt{Enc}_{can}$. These encoders play the same role as conventional bi-encoders in that condensing each query and candidate information into single vector embedding, but are trained separately from bi-encoders for the initial stage retriever.
$\mathbf{h}^{sent}_{q}$ $=\texttt{aggregator}(\texttt{Enc}_{qry}([\texttt{CLS}]\text{x}^{0}_{q}\ldots\text{x}^{k}_{q}))$ (3)
$\mathbf{h}^{sent}_{c_{q,j}}$ $=\texttt{aggregator}(\texttt{Enc}_{can}([\texttt{CLS}]\text{x}^{0}_{c_{q,j}}\ldots\text{x}^{k}_{c_{q,j}}))$ (4)
The aggregator function extracts [CLS] token embeddings from the last layer of each encoder [^2] For entity linking tasks, both the query (mention) and candidate (entity) sentences include custom special tokens that denote the locations of mention and entity words. These include [SEP], [query_start], [query_end], and [DOC] tokens following Wu et al. (2020)..
@@
The self-attention layer jointly processes concatenated embeddings of a query and all candidate sentences. This lightweight module enables parallel computation (efficient) and generates contextualized embeddings that capture query-candidate and candidate-candidate interaction (effective). Representing candidates together with self-attention enables fine-grained comparison among candidates. The self-attention layer consists of two layers of vanilla transformer encoder Vaswani et al. (2017) in Pytorch without positional encoding.
$\begin{split}\left[\mathbf{h}^{\texttt{CMC}}_{q};\mathbf{h}^{\texttt{CMC}}_{c_{q,1}};\ldots;\mathbf{h}^{\texttt{CMC}}_{c_{q,K}}\right]\\ =\texttt{SelfAttn}\Big{(}&\left[\mathbf{h}^{sent}_{q};\mathbf{h}^{sent}_{c_{q,1}};\ldots;\mathbf{h}^{sent}_{c_{q,K}}\right]\Big{)}\end{split}$ (5)
Consequently, CMC computes the final prediction $\hat{c}_{q}$ via dot products of query and candidate embeddings from the self-attention layer:
$\hat{c}_{q}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}\mathbf{h}^{\texttt{CMC}}_{q}\cdot\left(\mathbf{h}^{\texttt{CMC}}_{c_{q,j}}\right)^{\top}$ (6)

### v2

@@ Abstract
A common retrieve-and-rerank paradigm involves retrieving relevant candidates from a broad set using a fast bi-encoder (BE), followed by applying expensive but accurate cross-encoders (CE) to a limited candidate set. However, relying on this small subset is often susceptible to error propagation from the bi-encoders, which limits the overall performance. To address these issues, we propose the Comparing Multiple Candidates (CMC) framework. CMC compares a query and multiple embeddings of similar candidates (i.e., neighbors) through shallow self-attention layers, delivering rich representations contextualized to each other. Furthermore, CMC is scalable enough to handle multiple comparisons simultaneously. For example, comparing  10K candidates with CMC takes a similar amount of time as comparing 16 candidates with CE. Experimental results on the ZeSHEL dataset demonstrate that CMC, when plugged in between bi-encoders and cross-encoders as a seamless intermediate reranker (BE-CMC-CE), can effectively improve recall@k (+4.8%-p, +3.5%-p for R@16, R@64) compared to using only bi-encoders (BE-CE), with negligible slowdown (<7%). Additionally, to verify CMC’s effectiveness as the final-stage reranker in improving top-1 accuracy, we conduct experiments on downstream tasks such as entity, passage, and dialogue ranking. The results indicate that CMC is not only faster (11x) but also often more effective than cross-encoders  with improved prediction accuracy in Wikipedia entity linking (+0.7%-p) and DSTC7 dialogue ranking (+3.3%-p).
@@ 1 Introduction
Figure 1: Model architectures for retrieval tasks. (a), (b), and (c) are existing architectures. (d) is our proposed ‘Comparing Multiple Candidates (CMC)’ architecture, which computes compatibility score by comparing the embeddings of a query and K multiple candidates via self-attention layers. Contrary to (a)-(c), CMC can process multiple candidates at once rather than conducting several forward passes for each (query, candidate) pair.
@@
The two-stage approach of retrieval and reranking has become a predominant method in tasks such as entity linking (EL) Wu et al. (2020); Zhang and Stratos (2021); Xu et al. (2023), open-domain question answering (ODQA) Nogueira and Cho (2019); Agarwal et al. (2022b); Shen et al. (2022); Qu et al. (2020), and dialogue systems Mele et al. (2020). Typically, bi-encoders (BE) are used to efficiently retrieve relevant candidates from a large set of documents (e.g., knowledge base), and then cross-encoders (CE) effectively rerank only a limited subset of candidates already retrieved by BE (Nogueira and Cho (2019); Figure 1.a-b).
The current BE-CE approach, although widely adopted, has an efficiency-effectiveness trade-off and is susceptible to error propagation. When less accurate BE retrieves candidates, the whole framework risks the error propagation of missing the gold candidates due to inaccuracies from the retriever. Simply increasing the number of candidates is not a viable solution considering the slow serving time of CE [^1] For the serving time of cross-encoders, see §D.1.222Furthermore, increasing the number of candidates for CE does not necessarily improve end-to-end accuracy Wu et al. (2020). We confirm this in the experiments. See appendix D.6.. Consequently, users are faced with the dilemma of deciding which is worse: error propagation from BE versus the slow runtime of CE.
@@
To resolve this issue, various strategies have been proposed to find an optimal balance in the efficiency-effectiveness trade-off. Prior works (Khattab and Zaharia (2020); Zhang and Stratos (2021); Cao et al. (2020); Humeau et al. (2019)) have enhanced bi-encoder architectures with a late interaction component. However, these models only focus on single query-candidate pair interaction. Also, they sometimes require storing entire token embeddings per candidate sentence which results in tremendous memory use (Figure 1.c).
Our proposed Comparing Multiple Candidates (CMC) makes reranking easy by comparing similar candidates (i.e., neighbors) together. By jointly contextualizing the single vector embeddings from each candidate through shallow bi-directional self-attention layers, CMC achieves high prediction accuracy and runtime efficiency that are comparable to, or better, than existing methods which require single or multiple vector embeddings.
@@
In other words, CMC only takes a single forward pass for input $(\text{query},\text{candidate}_{1},...,\text{candidate}_{k})$ with a pre-computed single vector embedding. In contrast, models such as CE and other late interaction models take $k$ separate forward passes for input pairs $(\text{query},\text{candidate}_{1}),...,(\text{query},\text{candidate}_{k})$, sometimes requiring multiple vector embeddings per each candidate. CMC maintains both the efficiency of BE with pre-computed single-vector candidate embeddings, and the effectiveness of CE with interactions between query and multiple candidates (Figure 1.d).
Practitioners can plug in CMC as the seamless intermediate reranker (BE-CMC-CE) which can enhance retrieval performance with negligible extra latency. This improvement is crucial for preventing error propagation from the retrieval process, resulting in more reliable candidates for the final stage (Figure 2-3). On the other hand, CMC also can serve as a fast and effective final-stage reranker improving top-1 accuracy (BE-CMC). If there’s a time constraint, using CMC as the final reranker can be a good option, as running a cross-encoder requires significantly more time (Table 3; Figure 4).
@@
In experiments, we evaluate CMC on Zero-SHot Entity-Linking dataset (ZeSHEL; Logeswaran et al. (2019)) to investigate how much CMC seamlessly enhances a retriever’s performance when plugged in to BE (BE-CMC). The results show CMC provides higher recall than baseline retrievers at a marginal increase in latency (+0.07x; Table 7). Compared to standard BE-CE, plugging in CMC as the seamless intermediate reranker (BE-CMC-CE) can provide fewer, higher-quality candidates to CE, ultimately improving the accuracy of CE reranking. (Table 12). To examine the effectiveness of CMC which acts as the final stage reranker, we evaluate CMC on entity, passage, and dialogue ranking tasks. We observe that CMC outperforms CE on Wikipedia entity linking datasets (+0.7p accuracy) and DSTC7 dialogue ranking datasets (+3.3p MRR), requiring only a small amount (0.09x) of CE’s latency (Table 3).
The main contributions of the paper are as follows:
  • We present a novel reranker, CMC, which improves both accuracy and scalability. CMC contextualizes candidate representations with similar candidates (i.e., neighbors), instead of solely focusing on a single query-candidate pair (§3).
  • CMC can serve as the seamless intermediate reranker which can significantly improve retrieval performance with only a negligible increase in latency. This results in a more confident set of candidates for the final-stage reranker that improves end-to-end accuracy compared to conventional bi-encoders (§4.3)
@@
  • Experimental results show that the final stage reranking of CMC is highly effective on passage, entity, and dialogue ranking tasks compared to various baselines among the low-latency models (§4.4).
@@
  • Additionally, we show that CMC can benefit from domain transfer from sentence encoders while BE and many others cannot (§4.5).
@@
Figure 2: Overview of the proposed CMC framework that compares multiple candidates at once. CMC can seamlessly enhance retriever, finding top-K’ candidates, or function as a direct reranker which outputs top-1 candidate. Candidate embeddings for bi-encoders and CMC are both precomputed while query embeddings for bi-encoders and CMC are computed in parallel on the fly. After bi-encoders retrieve top-$K$ candidates, CMC indexes the corresponding candidate embeddings and passes through a two-layer transformer encoder. Here, the additional latency is limited to the execution of self-attention layers.
@@ 2.1 Retrieve and Rerank
Two-stage retrieval systems commonly consist of a fast retriever and a slow but accurate reranker. Although the retriever is fast, its top-1 accuracy tends to be suboptimal. Therefore, a candidate set $C_{q}=\{c_{q,1},c_{q,2},\ldots,c_{q,K}\}\subseteq\mathcal{C}$ whose elements are $K$ most relevant candidates in the corpus $\mathcal{C}$ is retrieved for further reranking.
A reranker $s_{\theta}(q,c_{q,j})(1\leq j\leq K)$ is a model trained to assign a fine-grained score between the query $q$ and each candidate $c_{q,j}$ from the relatively small set of candidates $C_{q}$. It is an expressive model that is slower but more accurate than the retriever. The candidate with the highest score $c_{q}^{*}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}s_{\theta}(q,c_{q,j})$ is the final output of the retrieve-and-rerank pipeline where query $q$ should be linked.
@@ 2.2 Related Work
In two-stage retrieval, the compatibility score between the query and candidate can be computed by diverse functions. Nogueira et al. (2019a) retrieve candidates using the bag-of-words BM25 retriever and then apply a cross-encoder reranker, transformer encoders that take the concatenated query and candidate tokens as input Logeswaran et al. (2019); Wu et al. (2020). Instead of BM25 retriever, other works Lee et al. (2019); Gillick et al. (2019); Karpukhin et al. (2020) employ a pre-trained language model for a bi-encoders retriever to encode a query and a candidate separately and get the compatibility score. The scalability of bi-encoders as a retriever arises from the indexing of candidates and maximum inner-product search (MIPS); however, they tend to be less effective than cross-encoders as candidate representations do not reflect the query’s information (Figure 1.a-b). To enhance the performance of bi-encoders, follow-up works propose a task-specific fine-tuned model Gao and Callan (2022), injecting graph information Wu et al. (2023); Agarwal et al. (2022a), and multi-view text representations Ma et al. (2021); Liu et al. (2023).
@@
Late interaction models, which typically function as either a retriever or a reranker, enhance bi-encoder architectures with a late interaction component between the query and the candidate.
Poly-encoder Humeau et al. (2019) and MixEncoder Yang et al. (2023) represent query information through cross-attention with a candidate to compute the matching score. However, these models have overlooked the opportunity to explore the interaction among candidates.
Sum-of-Max Khattab and Zaharia (2020); Zhang and Stratos (2021) and DeFormer Cao et al. (2020) rely on maximum similarity operations or extra cross-encoder layers on top of bi-encoders. However, they lack scalability due to the need to pre-compute and save every token embedding per each candidate. [^3] For example, 3.2TB is required for storing $\sim$5M entity descriptions from Wikipedia, each with 128 tokens. In contrast, storing a single vector embedding per entity description for bi-encoders only requires 23GB. As a collection of documents continuously changes and grows, this storage requirement poses practical limitations on managing and updating the document indices.
CMC differs from these models in its enhanced scalability by comparing a single embedding for each candidate. This approach provides a deeper exploration of relational dynamics from interactions across multiple candidates while improving time and memory efficiency.
@@
CMC is not the first approach to compare a list of documents to enhance ranking performance Han et al. (2020); Zhang et al. (2022); Xu et al. (2023). These listwise ranking methods process cross-encoder logits for the list $(\text{query},\text{candidate}_{1},\ldots,\text{candidate}_{\text{K}})$ to rerank $K$ candidates from cross-encoders. However, these approaches lack scalability and efficiency due to reliance on cross-encoder representations.
Unlike previous listwise ranking models, we propose a method that employs representations from independent sentence encoders rather than cross-encoders. Boosting scalability with independent representations, CMC can seamlessly enhance retrievers by maintaining prediction accuracy.
@@ 3.1 Model Architecture
Comparing Multiple Candidates, CMC, employs shallow self-attention layers to capture both query-candidate and candidate-candidate interactions. Unlike other late interaction models which compute the compatibility scores by only considering a single query-candidate pair Khattab and Zaharia (2020); Humeau et al. (2019); Yang et al. (2023), CMC compares each candidate to the query and other candidates at the same time (Figure 1.(d)). The self-attention layer in CMC processes the concatenated representations of the query and multiple candidates, derived from the independent query and candidate encoders. In this way, CMC obtains enhanced representations of the query and every candidate by contextualizing them with each other. Also, this architecture is scalable to a large set of corpus by pre-computing and indexing candidate embeddings. For example, processing 2K candidates only takes twice as long as processing 100 (Figure 4).
@@
Prior to CMC, the first-stage retriever (e.g., bi-encoders) retrieves the candidate set with K elements $C_{q}=\{c_{q,1},...c_{q,K}\}$ for query $q$. CMC then obtains the aggregated encoder output (e.g., [CLS] token embedding) of query sentence tokens $\mathbf{h}^{sent}_{q}$ and candidate sentence tokens $\mathbf{h}^{sent}_{c_{q,j}}$ from the query encoder $\texttt{Enc}_{qry}$ and the candidate encoder $\texttt{Enc}_{can}$. These encoders play the same role as conventional bi-encoders by condensing each query and candidate information into a single vector embedding but are trained separately from the first-stage stage retriever.
$\mathbf{h}^{sent}_{q}$ $=\texttt{agg}(\texttt{Enc}_{qry}([\texttt{CLS}]\text{x}^{0}_{q}\ldots\text{x}^{k}_{q}))$ (1)
$\mathbf{h}^{sent}_{c_{q,j}}$ $=\texttt{agg}(\texttt{Enc}_{can}([\texttt{CLS}]\text{x}^{0}_{c_{q,j}}\ldots\text{x}^{k}_{c_{q,j}}))$ (2)
$\text{x}_{q}$ and $\text{x}_{c_{q,j}}$ are tokens of each query and candidate. The aggregator function agg extracts [CLS] embedding from the last layer of encoder [^4] For entity linking tasks, both the query (mention) and candidate (entity) sentences include custom special tokens that denote the locations of mention and entity words. These include [SEP], [query_start], [query_end], and [DOC] tokens following Wu et al. (2020)..
@@
The shallow self-attention layers process concatenated embeddings of a query and all candidates. This lightweight module enables parallel computation (efficient) and outputs contextualized embeddings via interactions between query and candidates (effective). In the reranking perspective, Representing candidates together with self-attention layers (Attn) enables fine-grained comparison among candidates. The self-attention layers consist of two layers of vanilla transformer encoder Vaswani et al. (2017) in Pytorch without positional encoding.
$\left[\mathbf{h}^{\texttt{CMC}}_{q};\mathbf{h}^{\texttt{CMC}}_{c_{q,1}};\ldots;\mathbf{h}^{\texttt{CMC}}_{c_{q,K}}\right]=\texttt{Attn}\left(\left[\mathbf{h}^{sent}_{q};\mathbf{h}^{sent}_{c_{q,1}};\ldots;\mathbf{h}^{sent}_{c_{q,K}}\right]\right)$
Subsequently, the reranker computes the final prediction $c^{*}_{q}$ via dot products of query and candidate embeddings from the self-attention layer:
$c^{*}_{q}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}\mathbf{h}^{\texttt{CMC}}_{q}\cdot\left(\mathbf{h}^{\texttt{CMC}}_{c_{q,j}}\right)^{\top}$ (4)
3.2 Training
Optimization
The training objective is minimizing the cross-entropy loss regularized by the Kullback-Leibler (KL) divergence between the score distribution of the trained model and the bi-encoder. The loss function is formulated as:
$\mathcal{L}(q,\tilde{C}_{q})=\sum_{i=1}^{K}(-\lambda_{1}y_{i}\log(p_{i})+\lambda_{2}p_{i}\log\left(\frac{p_{i}}{r_{i}}\right))$
$y_{i}$ and $p_{i}$ are the ground truth and predicted probability for i-th candidate. The retriever’s probability for the candidate is represented as $r_{i}$. $\lambda_{1}$ and $\lambda_{2}$ are weights combining the two losses.
Negative Sampling
We sample hard negatives based on the first-stage retriever’s score for each query-candidate pair $(q,c_{q,j})$: $\forall j\in\{1,\ldots,K\}\setminus\{\text{gold index}\}$,
$\quad{c}_{q,j}\sim\frac{\exp(s_{\text{retriever}}(q,c_{q,j}))}{\sum_{\begin{subarray}{c}k=1\&\\ k\neq\text{gold index}\end{subarray}}^{K}\exp(s_{\text{retriever}}(q,c_{q,k}))}$ (6)
In experiments, CMC and other baselines follow the same optimization and negative sampling strategy. [^5] The code and link to datasets are available at https://github.com/yc-song/cmc